# Conditioning diffusion models by explicit forward-backward bridging

### v1

@@ Abstract
Given an unconditional diffusion model $\pi(x,y)$, using it to perform conditional simulation $\pi(x\mid y)$ is still largely an open question and is typically achieved by learning conditional drifts to the denoising SDE after the fact. In this work, we express conditional simulation as an inference problem on an augmented space corresponding to a partial SDE bridge. This perspective allows us to implement efficient and principled particle Gibbs and pseudo-marginal samplers marginally targeting the conditional distribution $\pi(x\mid y)$. Contrary to existing methodology, our methods do not introduce any additional approximation to the unconditional diffusion model aside from the Monte Carlo error. We showcase the benefits and drawbacks of our approach on a series of synthetic and real data examples.
@@ 1 Introduction
Denoising diffusion models (Song et al., 2021; Ho et al., 2020) have recently received a lot of attention as general-purpose samplers for generative modelling in diverse fields such as image analysis (Luo et al., 2023), protein folding (Wu et al., 2023), and statistical inference (Vargas et al., 2023). At the core, given a target distribution $\pi_{0}$ that we want to generate samples from, and a time-homogeneous stochastic differential equation (SDE, see, e.g., Karatzas and Shreve, 1991)
@@
with stationary distribution $\pi_{\mathrm{ref}}$, they sample from $\pi_{0}$ by “denoising” (1) via Doob’s $h$-transform (Rogers and Williams, 2000). Formally, let $\pi_{t}$ be the distribution of $X_{t}$ under (1), and assume that we know how to sample from $\pi_{T}$ for some $T>0$. We can obtain approximate samples from $\pi_{0}$ by sampling from
@@
where $B$ is another Brownian motion, and we write $f_{\mathrm{rev}}(U_{t},t)\coloneqq-f(U_{t})+\nabla\log\pi_{T-t}(U_{t})$. When $T\gg 1$, and under ergodicity guarantees (Meyn and Tweedie, 2009), $\pi_{T}\approx\pi_{\mathrm{ref}}$. Consequently, if (1) is chosen such that $\pi_{\mathrm{ref}}$ is easy to sample from, then the only remaining difficulty is computing the score $\nabla\log\pi_{T-t}$, which is typically overcome by training a neural network to approximate the score function (or drift) via score matching (Hyvärinen, 2005).
Once the score function has been computed, approximate samples from $\pi$ can be obtained by running a discretised simulation of (2)
$\begin{split}U_{k}&=U_{k-1}-(t_{k}-t_{k-1})\,f_{\mathrm{rev}}(U_{k-1},t_{k-1})+\sqrt{t_{k}-t_{k-1}}\,Z_{k},\\ U_{0}&\sim\pi_{\mathrm{ref}},\quad Z_{k}\sim\mathcal{N}(0,I_{d}),\end{split}$ (3)
for an integration grid $0=t_{0}<t_{1}<\cdots<t_{K}=T$, where $I_{d}$ is a $d$-dimensional diagonal matrix.
A typical choice for the forward SDE (1) is a Langevin diffusion targeting a standard Gaussian distribution, namely taking $f(x)=-x$, also known as the Ornstein–Uhlenbeck process. Variants of this choice have been proposed, such as the use of time-dependent diffusion coefficient (Song and Ermon, 2020) or non-Markovian dynamics (Song et al., 2020). A remarkable extension is the use of Schrödinger bridges (De Bortoli et al., 2021), which generalise the approach by forming SDE bridges between two arbitrary distributions, thereby allowing for the generation of samples from $\pi_{0}$ given samples from $\pi_{T}=\pi_{\mathrm{ref}}$, in finite time rather than asymptotically as $T\gg 1$ at the cost of learning both the forward and backward dynamics.
@@ 1.1 Problem formulation
Despite their success, using diffusion models to perform exact conditional simulation is still largely an open question, whereby the goal is to sample from the conditional distribution $\pi(x\mid y)$. Here, $x$ represents the sample of interest, unconditionally obtained from the diffusion model (3) and $y$ is a conditioning variable. In this article, we consider the problem of exact conditioning within diffusion models. This is a question posed by Bayesian inference in general (Martin et al., 2023), also known as inpainting and super-resolution within the computer vision community.
@@ 1.2 Contributions
We propose a new Markov chain Monte Carlo (MCMC, Meyn and Tweedie, 2009) method for conditioning within diffusion models, to obtain samples from $\pi(x\mid y)$. Our method, which we call forward-backward bridging (FBB), is based on the following observation: conditioning a diffusion model can be written as a partial SDE bridge, constructed by running a preliminary forward simulation, noising the initial states $(X_{0},y)$ into $(X_{T},Y_{T})$, and then a backward particle filter corresponding to bridging back to the target distribution $\pi_{0}(x\mid y)$. This perspective was originally proposed in Trippe et al. (2023) and in the work, developed independently from ours, of Dou and Song (2024).
In contrast to both these works, here we take a more principled approach of treating the conditioning problem as an inference problem on the joint $p(x_{0},x_{T},y_{T}\mid Y_{0}=y)$, as well as all intermediary missing steps. We achieve this by leveraging the dynamic structure of diffusion models and alternating between a forward noising process and a conditional sequential Monte Carlo (CSMC, Andrieu et al., 2010) algorithm, thereby forming a Gibbs sampler (Geman and Geman, 1984) alternatively sampling from the joint and conditional distributions of the diffusion model. This presents the substantial benefits of correcting the sampler for lack of ergodicity (i.e., the fact of using finite number of particles, and that $T<\infty$ being finite create biases in Trippe et al., 2023; Dou and Song, 2024) as well as providing a way to treat Schrödinger bridge samplers (De Bortoli et al., 2021; Shi et al., 2023) which cannot be handled by existing methodology.
@@
    We develop forward-backward bridging (FBB), a new particle Gibbs (Andrieu et al., 2010) method for conditioning diffusion models.
@@
    In Section 3, we show how, when the noising process is tractable and separable, which is the case for the usual denoising diffusion models, the method can be adapted to be run in-place, at zero memory cost using a pseudo-marginal approach (Andrieu et al., 2010). This provides a principled and unbiased generalisation of Trippe et al. (2023).
@@ 2 Forward-backward conditioning of diffusions
Let us consider a target density $\pi(x,y)=\pi(y)\,\pi(x\mid y)$ for which we have a diffusion sampler, and a value $y$ for conditioning. We remark that in contrast to Shi et al. (2022); Wu et al. (2023); Song et al. (2021), we do not assume the capability to sample from or evaluate $\pi(y\mid x)$. This section describes how we sample from $\pi(x\mid y)$ given a diffusion sampler for $\pi(x,y)$ without any additional training.
@@ 2.1 An abstract Gibbs sampler for the conditional distribution
$\begin{split}\mathop{}\!\mathrm{d}Y_{t}=\mu^{Y}(X_{t},Y_{t},t)\mathop{}\!\mathrm{d}t+\sigma_{t}^{Y}\mathop{}\!\mathrm{d}W^{Y}_{t},\qquad\mathop{}\!\mathrm{d}X_{t}=\mu^{X}(X_{t},Y_{t},t)\mathop{}\!\mathrm{d}t+\sigma_{t}^{X}\mathop{}\!\mathrm{d}W^{X}_{t},\end{split}$ (4)
with $(X_{0},Y_{0})\sim\pi$ and such that the pair $(X_{T},Y_{T})$ is distributed according to any reference measure $\pi_{T}=\pi_{\mathrm{ref}}$. We denote the measure associated with the diffusion (4) by $\mathbb{F}$. This type of SDE is typically given by standard diffusion models (Song et al., 2021) or Schrödinger bridges (De Bortoli et al., 2021, see also Appendix A).
@@
$\begin{split}\mathop{}\!\mathrm{d}V_{t}=\mu^{V}(U_{t},V_{t},t)\mathop{}\!\mathrm{d}t+\sigma_{t}^{V}\mathop{}\!\mathrm{d}W^{V}_{t},\qquad\mathop{}\!\mathrm{d}U_{t}=\mu^{U}(U_{t},V_{t},t)\mathop{}\!\mathrm{d}t+\sigma_{t}^{U}\mathop{}\!\mathrm{d}W^{U}_{t},\end{split}$ (5)
such that if $(U_{0},V_{0})\sim\pi_{\mathrm{ref}}$ then $(U_{T},V_{T})\sim\pi$ is distributed according to the target measure $\pi(x,y)$. We denote the measure associated with the backward diffusion (5) by $\mathbb{B}$. Our conditional sampler construction is based on the following abstract Gibbs sampler (Geman and Geman, 1984), which corresponds to alternatively propagating (4) from $t=0$ to $t=T$, and then sampling $U_{T}$ conditionally on $U_{0}=X_{T}$, $V_{0}=Y_{T}$, and $V_{T}=y$, coming from the forwards simulation pass. Formally, suppose that, at a Gibbs step $j\geq 0$, the path sample $X_{0}^{j}$ is distributed according to $\pi(x\mid y)$ for a given $y$, and consider the following procedure: {mdframed}[style=enumFrame]
@@
Then, $X^{j+1}_{0}$ is distributed according to the target $\pi(x\mid y)$.
@@ 2.2 Discrete-time formulation
For this procedure to be practically implementable, we now assume that we have access to the discretised forward and backward SDEs, which are given marginally by
$\begin{split}Y_{t_{k+1}}\sim\mathbb{F}^{Y}_{t_{k}}(\cdot\mid Y_{t_{k}},X_{t_{k}}),\\ X_{t_{k+1}}\sim\mathbb{F}^{X}_{t_{k}}(\cdot\mid Y_{t_{k}},X_{t_{k}}),\end{split}\qquad\begin{split}V_{t^{\prime}_{k+1}}\sim\mathbb{B}^{V}_{t^{\prime}_{k}}(\cdot\mid V_{t^{\prime}_{k}},U_{t^{\prime}_{k}}),\\ U_{t^{\prime}_{k+1}}\sim\mathbb{B}^{U}_{t^{\prime}_{k}}(\cdot\mid V_{t^{\prime}_{k}},U_{t^{\prime}_{k}}),\end{split}$ (6)
@@
This structure of the conditional sampler was first introduced in Trippe et al. (2023, albeit not explicitly as a Gibbs sampler) for the special case when the drift in (1) is separable, namely when the drifts $\mu^{X}$ and $\mu^{Y}$ in (1) do not depend on both $Y$ and $X$, respectively (we come back to this special case in Section 3).
The simulation in step 1 is directly implementable. On the other hand, step 2 a well-studied problem in the literature on Bayesian filtering (see, e.g., Särkkä and Svensson, 2023), and approximate solutions are available, for example, in the form of particle filters (for a recent review of these see Chopin and Papaspiliopoulos, 2020) which were employed in Trippe et al. (2023). However, for a finite number of Monte Carlo samples used, these methods provide biased (despite consistency) samples from the filtering distribution, therefore resulting in a biased sampler for targeting $\pi(\cdot\mid y)$ in Trippe et al. (2023). Instead, we implement a Markov kernel $\mathcal{K}$ keeping the filtering posterior $\mathbb{B}(u_{1:K}\mid v_{0:K},u_{0})$, for given $v_{0:K}$ and $u_{0}$, invariant within the Gibbs sampler, resulting in the following structure: {mdframed}[style=enumFrame]
@@
A natural algorithm for this purpose is given by conditional sequential Monte Carlo (CSMC, Andrieu et al., 2010), which modifies the standard particle filter algorithm into a Markov kernel. CSMC is known to be fast converging under weak conditions and scales well to long time horizons (Andrieu et al., 2018) as well as diminishing discretisation steps in continuous systems (provided an adequate implementation is used, see, e.g., Karppinen et al., 2023), including for as little as $N>2$ Monte Carlo samples (Lee et al., 2020). We assume that the conditional transition density $b^{V}_{k}$ of $\mathbb{B}_{k}$ can be evaluated, and we provide pseudo-code for the CSMC algorithm (as well as the unconditional version, which we use later in Section 3) applied to step 2 above in Algorithm 1. For more details on both the unconditional and conditional versions of the algorithm, we refer to Chopin and Papaspiliopoulos (2020, Chap. 10 and 15, respectively) and Andrieu et al. (2010). In practice, to increase statistical efficiency, we use lower-variance resampling schemes and a final index selection step as in Karppinen et al. (2023); Chopin and Singh (2015).
Algorithm 1 Conditional/unconditional particle filter for $\mathcal{K}(\cdot\mid x^{j}_{0:K})$.
0:  The current Markov chain state $x^{j}_{0:K}$, measurement path $y^{j}_{0:K}$, number of particles $N>1$, and conditional flag $F\in\{0,1\}$.
@@
1:  Sample $U^{n}_{1}\sim\mathbb{B}_{0}(\cdot\mid y^{j}_{K},x^{j}_{K})$ for $n\in\{1,\ldots,N\}$, and set $Z^{j}=0$
@@
4:  end if
5:  Set $w^{n}_{1}=b^{V}_{1}(y^{j}_{K-2}\mid y^{j}_{K-1},U^{n}_{1})$ for $n\in\{1,\ldots,N\}$ and set $\ell_{1}=\sum_{n=1}^{N}w_{1}^{n}$
6:  Normalise $w^{n}_{1}=w^{n}_{1}/\ell_{1}$ for $n\in\{1,\ldots,N\}$, and $Z^{j}=\frac{1}{N}\times Z\times\ell_{1}$
7:  for $k=2,\ldots,K$ do
8:     Sample $A^{n}_{k}=i$ with probability $w^{i}_{k-1}$ for $n\in\{1,\ldots,N\}$
9:     Sample $U_{k}^{n}\sim\mathbb{B}_{k}(\cdot\mid y^{j}_{K-k},U^{A^{n}_{k-1}}_{k-1})$ for $n\in\{1,\ldots,N\}$
10:     if $F=1$ then
11:        Sample $n_{k}^{*}\sim\mathcal{U}[1,N]$, and set $U^{n^{*}_{k}}_{k}=y^{j}_{k}$ and $A^{n^{*}_{k}}_{k}=n^{*}_{k-1}$
12:     end if
13:     Set $w^{n}_{k}=b^{v}_{k}(y^{j}_{K-k-1}\mid y^{j}_{K-k},U^{n}_{k})$ for $n\in\{1,\ldots,N\}$ and set $\ell_{k}=\sum_{n=1}^{N}w_{k}^{n}$
14:     Normalise $w^{n}_{k}=w^{n}_{k}/\ell_{k}$ for $n\in\{1,\ldots,N\}$, and $Z^{j}=\frac{1}{N}\times Z\times\ell_{k}$
15:  end for
16:  Sample $U_{K}^{n}\sim\mathbb{B}_{K-1}(\cdot\mid y^{j}_{K-1},U^{A^{n}_{K-1}}_{K-1})$ for $n\in\{1,\ldots,N\}$
17:  if $F=1$ then
18:     Sample $n_{k}^{*}\sim\mathcal{U}[1,N]$, and set $U^{n^{*}_{K}}_{K}=x_{0}^{j}$ and $A^{n^{*}_{k}}_{k}=n^{*}_{k-1}$
19:  end if
20:  Sample $B_{K}\in\{1,\ldots,N\}$ with probabilities $\{w^{n}_{K-1}\}_{n=1}^{N}$ and set $x^{j+1}_{0}=U^{B_{K}}_{K}$
Remark 2.1.

### v2

@@ Abstract
Given an unconditional diffusion model targeting a joint model $\pi(x,y)$, using it to perform conditional simulation $\pi(x\mid y)$ is still largely an open question and is typically achieved by learning conditional drifts to the denoising SDE after the fact. In this work, we express exact conditional simulation within the approximate diffusion model as an inference problem on an augmented space corresponding to a partial SDE bridge. This perspective allows us to implement efficient and principled particle Gibbs and pseudo-marginal samplers marginally targeting the conditional distribution $\pi(x\mid y)$. Contrary to existing methodology, our methods do not introduce any additional approximation to the unconditional diffusion model aside from the Monte Carlo error. We showcase the benefits and drawbacks of our approach on a series of synthetic and real data examples.
@@ 1 INTRODUCTION
Denoising diffusion models (Song et al.,, 2021; Ho et al.,, 2020) have recently received a lot of attention as general-purpose samplers for generative modelling in diverse fields such as image analysis (Luo et al.,, 2023), protein folding (Wu et al.,, 2023), and statistical inference (Vargas et al.,, 2023). At the core, given a target distribution $\pi_{0}$ that we want to generate samples from, and a time-homogeneous stochastic differential equation (SDE, see, e.g., Karatzas and Shreve,, 1991)
@@
with stationary distribution $\pi_{\mathrm{ref}}$, they sample from $\pi_{0}$ by “denoising” (1) via Doob’s $h$-transform (Rogers and Williams,, 2000). Formally, let $\pi_{t}$ be the distribution of $X_{t}$ under (1), and assume that we know how to sample from $\pi_{T}$ for some $T>0$. We can obtain approximate samples from $\pi_{0}$ by sampling from
@@
where $B$ is another Brownian motion, and we write $f_{\mathrm{rev}}(U_{t},t)\coloneqq-f(U_{t})+\nabla\log\pi_{T-t}(U_{t})$. When $T\gg 1$, and under ergodicity guarantees (Meyn and Tweedie,, 2009), $\pi_{T}\approx\pi_{\mathrm{ref}}$. Consequently, if (1) is chosen such that $\pi_{\mathrm{ref}}$ is easy to sample from, then the only remaining difficulty is computing the score $\nabla\log\pi_{T-t}$, which is typically overcome by training a neural network to approximate the score function (or drift) via score matching (Hyvärinen,, 2005).
Once the score function has been computed, approximate samples from $\pi$ can be obtained by running a discretised simulation of (2) as
$\begin{split}U_{k}&=U_{k-1}-\Delta_{k}\,f_{\mathrm{rev}}(U_{k-1},t_{k-1})+\sqrt{\Delta_{k}}\,Z_{k},\\ U_{0}&\sim\pi_{\mathrm{ref}},\quad Z_{k}\sim\mathcal{N}(0,I_{d}),\end{split}$ (3)
at any grid $0=t_{0}<\cdots<t_{K}=T$, where $I_{d}$ is a $d$-dimensional diagonal matrix, and $\Delta_{k}=t_{k}-t_{k-1}$.
A typical choice for the forward SDE (1) is a Langevin diffusion targeting a standard Gaussian distribution, namely taking $f(x)=-x$, also known as the Ornstein–Uhlenbeck process. Variants of this choice have been proposed, such as the use of time-dependent diffusion coefficient (Song and Ermon,, 2020) or non-Markovian dynamics (Song et al.,, 2020). A remarkable extension is the use of Schrödinger bridges (De Bortoli et al.,, 2021), which generalise the approach by forming SDE bridges between two arbitrary distributions, thereby allowing for the generation of samples from $\pi_{0}$ given samples from $\pi_{T}=\pi_{\mathrm{ref}}$, in finite time rather than asymptotically as $T\gg 1$ at the cost of learning both the forward and backward dynamics.
@@ 1.1 PROBLEM FORMULATION
Despite the success of diffusion models, performing exact conditional simulation within these is still largely an open question, whereby the goal is to sample from the conditional distribution $\pi(x\mid y)$ given a diffusion model representing the joint distribution $\pi(x,y)$. In this article, we assume that the model for $\pi(x,y)$ is correct and can be trusted so that we typically ignore unconditional modelling errors. We therefore consider the problem of exact conditioning within diffusion models. This is a question posed by Bayesian inference in general (Martin et al.,, 2023), also known as inpainting and super-resolution within the computer vision community.
@@ 1.2 CONTRIBUTIONS
We propose a new Markov chain Monte Carlo (MCMC, Meyn and Tweedie,, 2009) method for conditioning within diffusion models, to obtain samples from $\pi(x\mid y)$. Our method, which we call forward-backward bridging (FBB), is based on the following observation: conditioning a diffusion model can be written as a partial SDE bridge, constructed by running a preliminary forward simulation, noising the initial states $(X_{0},y)$ into $(X_{T},Y_{T})$, and then a backward particle filter corresponding to bridging back to the target distribution $\pi_{0}(x\mid y)$. This perspective was originally proposed in Trippe et al., (2023) and in the work, concurrent to ours, of Dou and Song, (2024)
In contrast to both these works, which approximate the conditional sampling procedure by ignoring errors arising from the use of a backward particle filter, here we take a more principled approach of treating the conditioning problem as an inference problem on the joint $p(x_{0},x_{T},y_{T}\mid Y_{0}=y)$, as well as all intermediary missing steps. We achieve this by leveraging the dynamic structure of diffusion models and alternating between a forward noising process and a conditional sequential Monte Carlo (CSMC, Andrieu et al.,, 2010) algorithm, thereby forming a Gibbs sampler (Geman and Geman,, 1984) alternatively sampling from the joint forward and conditional backward distributions of the diffusion model. This presents the substantial benefits of correcting the sampler for lack of ergodicity (i.e., the fact of using finite number of particles, and that $T<\infty$ being finite create biases in Trippe et al.,, 2023; Dou and Song,, 2024) as well as providing a way to treat Schrödinger bridge samplers (De Bortoli et al.,, 2021; Shi et al.,, 2023) which cannot be handled by existing methodology.
@@
    We develop forward-backward bridging (FBB), a new particle Gibbs (Andrieu et al.,, 2010) method for conditioning diffusion models.
@@
    In Section 3, we show how, when the noising process is tractable and separable, which is the case for the usual denoising diffusion models, the method can be adapted to be run in-place, at zero memory cost using a pseudo-marginal approach (Andrieu et al.,, 2010). This provides a principled and unbiased generalisation of Trippe et al., (2023).
@@ 2 FORWARD-BACKWARD CONDITIONING OF DIFFUSIONS
Let us consider a target density $\pi(x,y)=\pi(y)\,\pi(x\mid y)$ for which we have a diffusion sampler, and a value $y$ for conditioning. We remark that in contrast to Shi et al., (2022); Wu et al., (2023); Song et al., (2021), we do not assume the capability to sample from or evaluate $\pi(y\mid x)$. This section describes how we sample from $\pi(x\mid y)$ given a diffusion sampler for $\pi(x,y)$ without any additional training.
@@ 2.1 AN ABSTRACT GIBBS SAMPLER FOR THE CONDITIONAL DISTRIBUTION
$\begin{split}\mathop{}\!\mathrm{d}Y_{t}&=\mu^{Y}(X_{t},Y_{t},t)\mathop{}\!\mathrm{d}t+\sigma_{t}\mathop{}\!\mathrm{d}W^{Y}_{t},\\ \mathop{}\!\mathrm{d}X_{t}&=\mu^{X}(X_{t},Y_{t},t)\mathop{}\!\mathrm{d}t+\sigma_{t}\mathop{}\!\mathrm{d}W^{X}_{t},\end{split}$ (4)
with $(X_{0},Y_{0})\sim\pi$ and such that the pair $(X_{T},Y_{T})$ is distributed according to any reference measure $\pi_{T}=\pi_{\mathrm{ref}}$. We denote the measure associated with the diffusion (4) by $\mathbb{F}$, and we simplify using the same dispersion $\sigma_{t}$ although we are not limited to this setting. This type of SDE is typically given by standard diffusion models (Song et al.,, 2021) or Schrödinger bridges (De Bortoli et al.,, 2021, see also Appendix A).
@@
$\begin{split}\mathop{}\!\mathrm{d}V_{t}&=\mu^{V}(U_{t},V_{t},t)\mathop{}\!\mathrm{d}t+\sigma_{t}\mathop{}\!\mathrm{d}W^{V}_{t},\\ \mathop{}\!\mathrm{d}U_{t}&=\mu^{U}(U_{t},V_{t},t)\mathop{}\!\mathrm{d}t+\sigma_{t}\mathop{}\!\mathrm{d}W^{U}_{t},\end{split}$ (5)
such that if $(U_{0},V_{0})\sim\pi_{\mathrm{ref}}$ then $(U_{T},V_{T})\sim\pi$ is distributed according to the target measure $\pi(x,y)$. We denote the measure associated with the backward diffusion (5) by $\mathbb{B}$. Our conditional sampler construction is based on the following abstract Gibbs sampler (Geman and Geman,, 1984), which corresponds to alternatively propagating (4) from $t=0$ to $t=T$, and then sampling $U_{T}$ conditionally on $U_{0}=X_{T}$, $V_{0}=Y_{T}$, and $V_{T}=y$, coming from the forwards simulation pass. Formally, suppose that, at a Gibbs step $j\geq 0$, the path sample $X_{0}^{j}$ is distributed according to $\pi(x\mid y)$ for a given $y$, and consider the following procedure: {mdframed}[style=enumFrame]
@@
Then, $X^{j+1}_{0}$ distributes according to the target $\pi(x\mid y)$.
@@ 2.2 DISCRETE-TIME FORMULATION
For this procedure to be practically implementable, we now assume that we have access to the discretised forward and backward SDEs, given by
$\begin{split}Y_{t_{k+1}}\sim\mathbb{F}^{Y}_{t_{k}}(\cdot\mid Y_{t_{k}},X_{t_{k}}),\\ X_{t_{k+1}}\sim\mathbb{F}^{X}_{t_{k}}(\cdot\mid Y_{t_{k}},X_{t_{k}}),\end{split}\quad\begin{split}V_{t^{\prime}_{k+1}}\sim\mathbb{B}^{V}_{t^{\prime}_{k}}(\cdot\mid V_{t^{\prime}_{k}},U_{t^{\prime}_{k}}),\\ U_{t^{\prime}_{k+1}}\sim\mathbb{B}^{U}_{t^{\prime}_{k}}(\cdot\mid V_{t^{\prime}_{k}},U_{t^{\prime}_{k}}),\end{split}$ (6)
@@
This structure of the conditional sampler was first introduced in Trippe et al., (2023, albeit not explicitly as a Gibbs sampler) for the special case when the drift in (1) is separable, namely when the drifts $\mu^{X}$ and $\mu^{Y}$ in (1) do not depend on both $Y$ and $X$, respectively (we come back to this special case in Section 3).
The simulation in step 1 is directly implementable. On the other hand, step 2 a well-studied problem in the literature on Bayesian filtering (see, e.g., Särkkä and Svensson,, 2023), and approximate solutions are available, for example, in the form of particle filters (for a recent review of these see Chopin and Papaspiliopoulos,, 2020) which were employed in Trippe et al., (2023). However, for a finite number of Monte Carlo samples used, these methods provide biased (despite consistency) samples from the filtering distribution, therefore resulting in a biased sampler for targeting $\pi(\cdot\mid y)$ in Trippe et al., (2023). Instead, we implement a Markov kernel $\mathcal{K}$ keeping the filtering posterior $\mathbb{B}(u_{1:K}\mid v_{0:K},u_{0})$, for given $v_{0:K}$ and $u_{0}$, invariant within the Gibbs sampler, resulting in the following structure: {mdframed}[style=enumFrame]
@@
Remark 2.1.
In practice, an additional statistical bias remains. Indeed, due to $\mathbb{B}$ being a learned model (as well as $\mathbb{F}$ when treating of Schrödinger models), sampling from $\mathbb{F}(\cdot\mid x_{0}=X^{j}_{0},y_{0}=y)$ is not equivalent to sampling from $\mathbb{B}(\cdot\mid x_{0}=X^{j}_{0},y_{0}=y)$. Thankfully, this can be corrected too, and we detail the procedure in Appendix D.
A natural algorithm for this purpose is given by conditional sequential Monte Carlo (CSMC, Andrieu et al.,, 2010), which modifies the standard particle filter algorithm into a Markov kernel. CSMC is known to be fast converging under weak conditions and scales well to long time horizons (Andrieu et al.,, 2018) as well as diminishing discretisation steps in continuous systems (provided an adequate implementation is used, see, e.g., Karppinen et al.,, 2023), including for as little as $N>2$ Monte Carlo samples (Lee et al.,, 2020). We assume that the conditional transition density $b^{V}_{k}$ of $\mathbb{B}_{k}$ can be evaluated, and we provide pseudo-code for the CSMC algorithm (as well as the unconditional version, which we use later in Section 3) applied to step 2 above in Algorithm 1. For more details of the algorithm, we refer to Chopin and Papaspiliopoulos, (2020, Chaps. 10 and 15) and Andrieu et al., (2010). In practice, to increase statistical efficiency, we use lower-variance resampling schemes and a final index selection step as in Karppinen et al., (2023); Chopin and Singh, (2015).
Remark 2.2.
Remark 2.3.
The computational complexity of the method is exactly that of a particle filter using the same number of particles, and we therefore do not pay any additional cost compared to Trippe et al., (2023) when running the method for a single iteration of our Markov chain. On the other hand, the memory cost would generally scale as $O(T\times N)$, due to the need to have access to $x^{j}_{0:K}$. In Section 3.1, we discuss how this memory cost can be fully mitigated when the forward noising SDE is separable and Gaussian.
Algorithm 1 (Un)conditional particle filter.
0:  The current Markov chain state $x^{j}_{0:K}$, measurement path $y^{j}_{0:K}$, number of particles $N>1$, and conditional flag $F\in\{0,1\}$. Note that we shorten $n\!\upharpoonright_{1}^{N}$ as “$n=1,\ldots,N$”.
@@
1:  Sample $\{U^{n}_{1}\}_{n=1}^{N}\sim\mathbb{B}_{0}(\cdot\mid y^{j}_{K},x^{j}_{K})$, and set $Z^{j}=0$
@@
4:  Set $\{w^{n}_{1}=b^{V}_{1}(y^{j}_{K-2}\mid y^{j}_{K-1},U^{n}_{1})\}_{n=1}^{N}$ and $\ell_{1}=\sum_{n=1}^{N}w_{1}^{n}$
5:  Normalise $w^{n}_{1}=w^{n}_{1}/\ell_{1}$ for $n\!\upharpoonright_{1}^{N}$, and $Z^{j}=\frac{Z\,\ell_{1}}{N}$
6:  for $k=2,\ldots,K$ do
7:     Sample $A^{n}_{k}=i$ with probability $w^{i}_{k-1}$ for $n\!\upharpoonright_{1}^{N}$
8:     Sample $U_{k}^{n}\sim\mathbb{B}_{k}(\cdot\mid y^{j}_{K-k},U^{A^{n}_{k-1}}_{k-1})$ for $n\!\upharpoonright_{1}^{N}$
9:     if $F=1$ then
10:        Sample $n_{k}^{*}\sim\mathcal{U}[1,N]$, and set $U^{n^{*}_{k}}_{k}=y^{j}_{k}$ and $A^{n^{*}_{k}}_{k}=n^{*}_{k-1}$
11:     Set $\{w^{n}_{k}=b^{v}_{k}(y^{j}_{K-k-1}\mid y^{j}_{K-k},U^{n}_{k})\}_{n=1}^{N}$ and $\ell_{k}=\sum_{n=1}^{N}w_{k}^{n}$
12:     Normalise $w^{n}_{k}=w^{n}_{k}/\ell_{k}$ for $n\!\upharpoonright_{1}^{N}$, and $Z^{j}=\frac{Z\,\ell_{k}}{N}$
13:  Sample $U_{K}^{n}\sim\mathbb{B}_{K-1}(\cdot\mid y^{j}_{K-1},U^{A^{n}_{K-1}}_{K-1})$ for $n\!\upharpoonright_{1}^{N}$
14:  if $F=1$ then
15:     Sample $n_{k}^{*}\sim\mathcal{U}[1,N]$, and set $U^{n^{*}_{K}}_{K}=x_{0}^{j}$ and $A^{n^{*}_{k}}_{k}=n^{*}_{k-1}$
16:  Sample $B_{K}\in\{1,\ldots,N\}$ with probabilities $\{w^{n}_{K-1}\}_{n=1}^{N}$ and set $x^{j+1}_{0}=U^{B_{K}}_{K}$